OpenStudy (anonymous):

Please help me. Explain 6x-1=8x+5

OpenStudy (anonymous):

Hi! Well, you can solve this by moving all the "x" terms to one side, and all the plain numbers, also called constants, to the other side.

OpenStudy (anonymous):

to start you will subtract 6x and 5 from both sides and get this: -1 - 5 = 8x - 6x now simplify by combining like variables: -6 = 2x now divide both sides by 2 -6/2 = x now simplify: -3 = x

OpenStudy (anonymous):

combine the all unknown together and combine whole number too. put unknown in separate part like unknown at left hand side while whole number at right hand side. then solve like usual.

OpenStudy (anonymous):

but must remember to move any subject at other side the sign must be change like from negative to positive.

OpenStudy (anonymous):

\[6x-1=8x+5 ...... 6x-8x=1+5.... -2x=6...... -2x /-2=6/-2....... x=\]